OpenStudy (anonymous):

log10(2x + 6) + log10(x − 1) = 1 find real number solutions

OpenStudy (lgbasallote):

log_10(2x+6)(x-1) = 1 10 = (2x+6)(x-1) x = 2 x = 11/2

OpenStudy (anonymous):

it said the answe 2 and 11/2 are incoreect

OpenStudy (anonymous):

(1) change logarithmic function to exponential function. .